find the point of intersection for the lines below using the substitution
3x-4y=6
x=3y+2

Answer by stanbon(75887)

find the point of intersection for the lines below using the substitution
1st: 3x-4y=6
2nd: x=3y+2
----------
Substitute for "x" in 1st to get:
3(3y+2)-4y=6
9y+6-4y=6
5y=0
y=0
Then x=3(0)+2=2
SOLUTION:
x=2; y=0
Cheers,
Stan H.